A brachiosaurus weighed 80,000 kg, but it's brain weights only .2 kg. What would your brain weigh if you were 150 pounds if your ratio was the same as the brachiosaurus?
Answer by checkley71(8403)

.2/80,000=X/150 CROSS MULTIPLY.
80,000X=.2*150
80,000X=30
X=30/80,000
X=.000375 POUNDS WOULD BE THE WEIGHT OF THE BRAIN IN A 150 POUND PERSON.